# ARC: A Generalist Graph Anomaly Detector with In-Context Learning

### v1

@@ Abstract
Graph anomaly detection (GAD), which aims to identify abnormal nodes that differ from the majority within a graph, has garnered significant attention. However, current GAD methods necessitate training specific to each dataset, resulting in high training costs, substantial data requirements, and limited generalizability when being applied to new datasets and domains. To address these limitations, this paper proposes ARC, a generalist GAD approach that enables a “one-for-all” GAD model to detect anomalies across various graph datasets on-the-fly. Equipped with in-context learning, ARC can directly extract dataset-specific patterns from the target dataset using few-shot normal samples at the inference stage, without the need for retraining or fine-tuning on the target dataset. ARC comprises three components that are well-crafted for capturing universal graph anomaly patterns: 1) smoothness-based feature Alignment module that unifies the features of different datasets into a common and anomaly-sensitive space; 2) ego-neighbor Residual graph encoder that learns abnormality-related node embeddings; and 3) cross-attentive in-Context anomaly scoring module that predicts node abnormality by leveraging few-shot normal samples. Extensive experiments on multiple benchmark datasets from various domains demonstrate the superior anomaly detection performance, efficiency, and generalizability of ARC.
@@ 1 Introduction
Graph anomaly detection (GAD) aims to distinguish abnormal nodes that show significant dissimilarity from the majority of nodes in a graph. GAD has broad applications across various real-world scenarios, such as fraud detection in financial transaction networks [1] and rumor detection in social networks [2]. As a result, GAD has attracted increasing research attention in recent years [3, 4, 5, 6, 7]. Conventional GAD methods employ shallow mechanisms to model node-level abnormality [8, 9, 10]; however, they face limitations in handling high-dimensional features and complex interdependent relations on graphs. Recently, graph neural network (GNN)-based approaches have emerged as the go-to solution for the GAD problem due to their superior performance [4, 6]. Some GNN-based GAD approaches regard GAD as a supervised binary classification problem and use specifically designed GNN architectures to capture anomaly patterns [6, 11, 12, 13]. Another line of approaches targets the more challenging unsupervised paradigm, employing various unsupervised learning objectives and frameworks to identify anomalies without relying on labels [4, 14, 15, 16].
@@ 2 Related Work
Anomaly Detection. Anomaly detection (AD) aims to identify anomalous samples that deviate from the majority of samples [20]. Mainstream AD methods focus on unsupervised settings and employ various unsupervised techniques to build the models [21, 22, 23, 24, 25, 26]. To enhance the generalizability of AD methods across diverse datasets, RegAD [27] considers few-shot setting and trains a single generalizable model capable of being applied to new in-domain data without re-training or fine-tuning. WinCLIP [28] utilizes visual-language models (VLMs, e.g., CLIP [29]) with well-crafted text prompts to perform zero/few-shot AD for image data. InCTRL [17], as the first generalist AD approach, integrates in-context learning and VLMs to achieve domain-agnostic image AD with a single model. However, due to their heavy reliance on pre-trained vision encoders/VLMs and image-specific designs, these approaches excel in AD for image data but face challenges when applied to graph data.
Graph Anomaly Detection (GAD). In this paper, we focus on the node-level AD on graphs and refer to it as “graph anomaly detection (GAD)” following [6, 30, 31]. While shallow methods [8, 9, 10] show limitations in handling complex real-world graphs [4], the advanced approaches are mainly based on GNNs [32]. The GNN-based approaches can be divided into supervised and unsupervised approaches [3, 5, 7]. Supervised GAD approaches assume that the labels of both normal and anomalous nodes are available for model training [7]. Hence, related studies mainly introduce GAD methods in a binary classification paradigm [6, 11, 12, 13, 33, 34]. In contrast, unsupervised GAD approaches do not require any labels for model training. They employ several unsupervised learning techniques to learn anomaly patterns on graph data, including data reconstruction [4, 31, 35], contrastive learning [14, 36, 37], and other auxiliary objectives [15, 16, 38]. Nevertheless, all the above methods adhere to the conventional paradigm of “one model for one dataset”. Although some GAD approaches [39, 40] can handle cross-domain scenarios, their requirement for high correlation (e.g., aligned node features) between source and target datasets limits their generalizability. Differing from existing methods, our proposed ARC is a “one-for-all” GAD model capable of identifying anomalies across target datasets from diverse domains, without the need for re-training or fine-tuning.
@@ 4.1 Smoothness-Based Feature Alignment
From the perspective of graph signal processing, Tang et al. [6] have demonstrated that the inverse of the low-frequency energy ratio monotonically increases with the anomaly degree. In other words, high-frequency graph signals tend to play a more significant role in detecting anomalies. Similar findings have also been observed from the standpoint of spatial GNNs [34, 51], where heterophily information has been shown to be crucial in discriminating anomalies. Motivated by these findings, can we develop a metric to gauge the contribution of each feature to GAD based on its frequency/heterophily?
Considering its correlation to frequency [52] and heterophily [53], in this paper, we select feature-level smoothness as the measure for contribution. Formally, given a graph with a normalized feature matrix , the smoothness of the -th feature dimension is defined as:
@@ 4.2 Ego-Neighbor Residual Graph Encoder
To address the above issues, we design an ego-neighbor residual graph encoder for ARC. Equipped with a residual operation, the encoder can capture multi-hop affinity patterns of each node, providing valuable and comprehensive information for anomaly identification. Similar to the “propagation then transformation” GNN architecture in SGC [56], our graph encoder consists of three steps: multi-hop propagation, shared MLP-based transformation, and ego-neighbor residual operation. In the first two steps, we perform propagation on the aligned feature matrix for iterations, and then conduct transformation on the raw and propagated features with a shared MLP network:
where , is the propagated feature matrix at the -th iteration, is the transformed representation matrix at the -th iteration, and is the normalized adjacency matrix [54, 56]. Note that, unlike most GNNs that only consider the features/representations after -iter propagation, here we incorporate both the raw features and intermediate propagated features and transform them into the same representation space. After obtaining , we calculate the residual representations by taking the difference between () and , and then concatenate the multi-hop residual representations to form the final embeddings:
@@ 4.3 Cross-Attentive In-Context Anomaly Scoring
Discussion. The design of cross-attentive in-context anomaly scoring follows a basic assumption: normal query nodes have similar patterns to several context nodes, and hence their embeddings can be easily represented by the linear combination of context node embeddings. Consequently, given a normal node, its original and reconstructed embeddings can be close to each other in the embedding space. In contrast, abnormal nodes may display distinct patterns compared to normal ones, making it difficult to depict their corresponding abnormal query embeddings using context embeddings. As a result, their drift distance can be significantly larger. Fig. 4 provides examples for the scenarios of (a) single-class normal and (b) multi-class normal. In both cases, the drift distance (➞) can be a significant indicator for distinguishing anomaly (5) from normal nodes (1~4). Interestingly, if the attention matrix assigns uniform weights to all context nodes, then our scoring module becomes a one-class classification model [21]. This property ensures the anomaly detection capability of ARC even without extensive training. A detailed discussion is conducted in Appendix D.
@@ 5.2 Experimental Results
Performance Comparison. Table 1 shows the comparison results of ARC with baseline methods in terms of AUROC. Results in AUPRC are provided in Appendix G.1. We have the following observations. ❶ ARC demonstrates strong anomaly detection capability in the generalist GAD scenario, without any fine-tuning. Specifically, ARC achieves state-of-the-art performance on 5 out of 8 datasets and demonstrates competitive performance on the remainder.
On several datasets, ARC demonstrates significant improvement compared to the best baseline (e.g., 18.1% on Cora, 18.4% on CiteSeer, and 18.1% on Amazon). ❷ Simply pre-training the dataset-specific GAD methods typically results in poor generalization capability to new datasets. Specifically, the AUROC of the majority of “pre-train only” approaches is close to random guessing (50%) or even lower. ❸ With dataset-specific fine-tuning, the baseline methods achieve better performance in the majority of cases. However, the improvement can be minor or even negative in some cases, demonstrating the limitations of fine-tuning. Additionally, we observe that their performance is sometimes lower than the reported results from training models from scratch [4, 14, 16], indicating potential risk of negative transfer within the "pre-train & fine-tune" paradigm. ❹ Unsupervised baselines (except HCM-A) generally outperform the supervised ones, highlighting the difficulty of training a generalist GAD model using the binary classification paradigm.
@@ Appendix A Detailing Related Work
Anomaly Detection. The objective of anomaly detection (AD) is to identify anomalous samples that deviate from the majority of samples [20]. Due to the difficulty of collecting labeled anomaly data, mainstream AD methods mainly focus on unsupervised settings. To capture anomaly patterns without guidance by annotated labels, existing studies employ several advanced techniques to learn powerful AD models, such as one-class classification [21, 22], distance measurement [23, 62], data reconstruction [24, 63, 64], generative models [25, 65, 66], and self-supervised learning [26, 67]. For example, DeepSVDD [21] introduces a fully deep one-class classification objective for unsupervised anomaly detection, optimizing a data-enclosing hypersphere in output space to extract common factors of variation and demonstrating theoretical properties such as the -property. AnoDDPM [65], a simplex noise-based approach for anomaly detection, enhances anomaly capture with a partial diffusion strategy and multiscale simplex noise processing, facilitating faster inference and training on high-resolution images. While effective, these approaches are tailored to identify abnormal samples within a predetermined target dataset (i.e., the dataset for training), restricting their generalizability to new domains.
@@
Anomaly Detection on Graph Data. Based on the granularity of anomaly samples within graph data, existing AD approaches can be primarily categorized into three classes: node-level [4, 6], edge-level [72, 73], and graph-level [74, 75] AD. Due to its broad real-world applications, node-level AD receives the most research attention [3]. In this paper, we focus on the node-level AD and refer to it as “graph anomaly detection (GAD)”, following the convention of most previous papers [6, 30, 31].
Early GAD methods aimed to detect anomalies through shallow mechanisms. For example, AMEN [8] detects anomalies by utilizing the attribute correlations of nodes in each ego-network on the attribute network. In addition, residual analysis is another common method to measure the anomalies of nodes on an attribute network. In particular, Radar [9] characterizes the residuals of attribute information and their coherence with network information for anomaly detection. Further, ANOMALOUS [10] proposes joint learning of attribute selection and anomaly detection as a whole based on CUR decomposition and residual analysis. Despite the success of these methods on low-dimensional attribute network data, they do not work well when the network has a complex structure and high-dimensional attributes due to the limitations of their shallow mechanisms.
@@
In contrast, unsupervised GAD approaches do not require any labels for model training. Similar to unsupervised AD for image data, unsupervised GAD approaches employ several unsupervised learning techniques to learn anomaly patterns on graph data, including data reconstruction [4, 31, 35], contrastive learning [14, 36, 37], and other auxiliary objectives [15, 16, 38]. For example, DOMINANT [4] is a reconstruction-based approach that employs a graph convolution autoencoder to reconstruct both the adjacency and attribute matrices simultaneously, assessing node abnormality through a weighted sum of the reconstruction error terms. Similarly, ComGA [31] is a community-aware attribute GAD framework based on tailored GCNs to capture local, global, and structural anomalies. CoLA [14], the first contrastive self-supervised learning framework for GAD, samples novel contrast instance pairs in an unsupervised manner, utilizing contrastive learning to capture local information. HCM-A [15] integrates both local and global contextual information, employs hop count prediction as a self-supervised task, and utilizes Bayesian learning to enhance anomaly identification. TAM [16] optimizes the proposed anomaly metric (affinity) on the truncated graph end-to-end, considering one-class homophily and local affinity. Nevertheless, all the above methods adhere to the conventional paradigm of “one model for one dataset”.
@@
Very recently, few studies applied ICL to graph learning and GNNs. As an initial attempt to use ICL for GNNs, PRODIGY [47] leverages a prompt graph-based framework to conduct few-shot ICL for node-level classification and edge-level prediction tasks. Further, UniLP [48] introduces ICL to resolve conflicting connectivity patterns caused by distributional differences between different graphs. Nevertheless, these methods require context samples in multiple classes for ICL during inference. Then, how to leverage one-class context samples with ICL in the scenario of generalist GAD remains an open question.
@@ Appendix D Discussion of Cross-Attentive In-Context Anomaly Scoring
In this section, we first introduce the basic definition of one-class classification (OC) model, and then discuss the connection between one-class classification and cross-attentive in-context anomaly scoring module (C-AS for short).

### v2

@@ Abstract
Graph anomaly detection (GAD), which aims to identify abnormal nodes that differ from the majority within a graph, has garnered significant attention. However, current GAD methods necessitate training specific to each dataset, resulting in high training costs, substantial data requirements, and limited generalizability when being applied to new datasets and domains. To address these limitations, this paper proposes ARC, a generalist GAD approach that enables a “one-for-all” GAD model to detect anomalies across various graph datasets on-the-fly. Equipped with in-context learning, ARC can directly extract dataset-specific patterns from the target dataset using few-shot normal samples at the inference stage, without the need for retraining or fine-tuning on the target dataset. ARC comprises three components that are well-crafted for capturing universal graph anomaly patterns: 1) smoothness-based feature Alignment module that unifies the features of different datasets into a common and anomaly-sensitive space; 2) ego-neighbor Residual graph encoder that learns abnormality-related node embeddings; and 3) cross-attentive in-Context anomaly scoring module that predicts node abnormality by leveraging few-shot normal samples. Extensive experiments on multiple benchmark datasets from various domains demonstrate the superior anomaly detection performance, efficiency, and generalizability of ARC. The source code of ARC is available at https://github.com/yixinliu233/ARC.
@@ 1 Introduction
Graph anomaly detection (GAD) aims to distinguish abnormal nodes that show significant dissimilarity from the majority of nodes in a graph. GAD has broad applications across various real-world scenarios, such as fraud detection in financial transaction networks [1] and rumor detection in social networks [2]. As a result, GAD has attracted increasing research attention in recent years [3, 4, 5, 6, 7, 8]. Conventional GAD methods employ shallow mechanisms to model node-level abnormality [9, 10, 11]; however, they face limitations in handling high-dimensional features and complex interdependent relations on graphs. Recently, graph neural network (GNN)-based approaches have emerged as the go-to solution for the GAD problem due to their superior performance [4, 6]. Some GNN-based GAD approaches regard GAD as a supervised binary classification problem and use specifically designed GNN architectures to capture anomaly patterns [6, 12, 13, 14]. Another line of approaches targets the more challenging unsupervised paradigm, employing various unsupervised learning objectives and frameworks to identify anomalies without relying on labels [4, 15, 16, 17].
@@ 2 Related Work
Anomaly Detection. Anomaly detection (AD) aims to identify anomalous samples that deviate from the majority of samples [21]. Mainstream AD methods focus on unsupervised settings and employ various unsupervised techniques to build the models [22, 23, 24, 25, 26, 27, 28, 29]. To enhance the generalizability of AD methods across diverse datasets, RegAD [30] considers few-shot setting and trains a single generalizable model capable of being applied to new in-domain data without re-training or fine-tuning. WinCLIP [31] utilizes visual-language models (VLMs, e.g., CLIP [32]) with well-crafted text prompts to perform zero/few-shot AD for image data. InCTRL [18], as the first generalist AD approach, integrates in-context learning and VLMs to achieve domain-agnostic image AD with a single model. However, due to their heavy reliance on pre-trained vision encoders/VLMs and image-specific designs, these approaches excel in AD for image data but face challenges when applied to graph data.
Graph Anomaly Detection (GAD). In this paper, we focus on the node-level AD on graphs and refer to it as “graph anomaly detection (GAD)” following [6, 33, 34]. While shallow methods [9, 10, 11] show limitations in handling complex real-world graphs [4], the advanced approaches are mainly based on GNNs [35]. The GNN-based approaches can be divided into supervised and unsupervised approaches [3, 5, 7]. Supervised GAD approaches assume that the labels of both normal and anomalous nodes are available for model training [7]. Hence, related studies mainly introduce GAD methods in a binary classification paradigm [6, 12, 13, 14, 36, 37]. In contrast, unsupervised GAD approaches do not require any labels for model training. They employ several unsupervised learning techniques to learn anomaly patterns on graph data, including data reconstruction [4, 34, 38], contrastive learning [15, 39, 40], and other auxiliary objectives [16, 17, 41, 42]. Nevertheless, all the above methods adhere to the conventional paradigm of “one model for one dataset”. Although some GAD approaches [43, 44] can handle cross-domain scenarios, their requirement for high correlation (e.g., aligned node features) between source and target datasets limits their generalizability. Differing from existing methods, our proposed ARC is a “one-for-all” GAD model capable of identifying anomalies across target datasets from diverse domains, without the need for re-training or fine-tuning.
@@ 4.1 Smoothness-Based Feature Alignment
From the perspective of graph signal processing, Tang et al. [6] have demonstrated that the inverse of the low-frequency energy ratio monotonically increases with the anomaly degree. In other words, high-frequency graph signals tend to play a more significant role in detecting anomalies. Similar findings have also been observed from the standpoint of spatial GNNs [37, 55], where heterophily information has been shown to be crucial in discriminating anomalies. Motivated by these findings, can we develop a metric to gauge the contribution of each feature to GAD based on its frequency/heterophily? Considering its correlation to frequency [56] and heterophily [57, 58, 59, 60], in this paper, we select feature-level smoothness as the measure for contribution. Formally, given a graph with a normalized feature matrix , the smoothness of the -th feature dimension is defined as:
@@ 4.2 Ego-Neighbor Residual Graph Encoder
To address the above issues, we design an ego-neighbor residual graph encoder for ARC. Equipped with a residual operation, the encoder can capture multi-hop affinity patterns of each node, providing valuable and comprehensive information for anomaly identification. Similar to the “propagation then transformation” GNN architecture in SGC [63], our graph encoder consists of three steps: multi-hop propagation, shared MLP-based transformation, and ego-neighbor residual operation. In the first two steps, we perform propagation on the aligned feature matrix for iterations, and then conduct transformation on the initial and propagated features with a shared MLP network:
where , is the propagated feature matrix at the -th iteration, is the transformed representation matrix at the -th iteration, and is the normalized adjacency matrix [61, 63]. Note that, unlike most GNNs that only consider the features/representations after -iter propagation, here we incorporate both the initial features and intermediate propagated features and transform them into the same representation space. After obtaining , we calculate the residual representations by taking the difference between () and , and then concatenate the multi-hop residual representations to form the final embeddings:
@@ 4.3 Cross-Attentive In-Context Anomaly Scoring
Discussion. The design of cross-attentive in-context anomaly scoring follows a basic assumption: normal query nodes have similar patterns to several context nodes, and hence their embeddings can be easily represented by the linear combination of context node embeddings. Consequently, given a normal node, its original and reconstructed embeddings can be close to each other in the embedding space. In contrast, abnormal nodes may display distinct patterns compared to normal ones, making it difficult to depict their corresponding abnormal query embeddings using context embeddings. As a result, their drift distance can be significantly larger. Fig. 4 provides examples for the scenarios of (a) single-class normal and (b) multi-class normal111Specific definitions see Appendix D.1. In both cases, the drift distance (➞) can be a significant indicator for distinguishing anomaly (5) from normal nodes (1~4). Interestingly, if the attention matrix assigns uniform weights to all context nodes, then our scoring module becomes a one-class classification model [22]. This property ensures the anomaly detection capability of ARC even without extensive training. A detailed discussion is conducted in Appendix D.2.
@@ 5.1 Experimental Setup
5.2 Experimental Results
Performance Comparison. Table 1 shows the comparison results of ARC with baseline methods in terms of AUROC. Results in AUPRC are provided in Appendix G.1.
We have the following observations. ❶ ARC demonstrates strong anomaly detection capability in the generalist GAD scenario, without any fine-tuning. Specifically, ARC achieves state-of-the-art performance on 5 out of 8 datasets and demonstrates competitive performance on the remainder. On several datasets, ARC demonstrates significant improvement compared to the best baseline (e.g., 21.1% on Cora, 18.8% on CiteSeer, and 36.6% on Amazon). ❷ Simply pre-training the dataset-specific GAD methods typically results in poor generalization capability to new datasets. Specifically, the AUROC of the majority of “pre-train only” approaches is close to random guessing (50%) or even lower. ❸ With dataset-specific fine-tuning, the baseline methods achieve better performance in the majority of cases. However, the improvement can be minor or even negative in some cases, demonstrating the limitations of fine-tuning. Additionally, we observe that their performance is sometimes lower than the reported results from training models from scratch [4, 15, 17], indicating potential risk of negative transfer within the "pre-train & fine-tune" paradigm. ❹ Unsupervised baselines (except HCM-A) generally outperform the supervised ones, highlighting the difficulty of training a generalist GAD model using the binary classification paradigm.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This research was partly funded by Australian Research Council (ARC) under grants FT210100097 and DP240101547 and the CSIRO – National Science Foundation (US) AI Research Collaboration Program.
@@ Appendix A Detailing Related Work
Anomaly Detection. The objective of anomaly detection (AD) is to identify anomalous samples that deviate from the majority of samples [21]. Due to the difficulty of collecting labeled anomaly data, mainstream AD methods mainly focus on unsupervised settings. To capture anomaly patterns without guidance by annotated labels, existing studies employ several advanced techniques to learn powerful AD models, such as one-class classification [22, 23], distance measurement [24, 69], data reconstruction [25, 70, 71], generative models [26, 72, 73], and self-supervised learning [27, 74]. For example, DeepSVDD [22] introduces a fully deep one-class classification objective for unsupervised anomaly detection, optimizing a data-enclosing hypersphere in output space to extract common factors of variation and demonstrating theoretical properties such as the -property. AnoDDPM [72], a simplex noise-based approach for anomaly detection, enhances anomaly capture with a partial diffusion strategy [75] and multiscale simplex noise processing, facilitating faster inference and training on high-resolution images. While effective, these approaches are tailored to identify abnormal samples within a predetermined target dataset (i.e., the dataset for training), restricting their generalizability to new domains.
@@
Anomaly Detection on Graph Data. Based on the granularity of anomaly samples within graph data, existing AD approaches can be primarily categorized into three classes: node-level [4, 6], edge-level [80, 81, 82], and graph-level [83, 84, 85, 86, 87, 88] AD. Due to its broad real-world applications, node-level AD receives the most research attention [3]. In this paper, we focus on the node-level AD and refer to it as “graph anomaly detection (GAD)”, following the convention of most previous papers [6, 33, 34].
Early GAD methods aimed to detect anomalies through shallow mechanisms. For example, AMEN [9] detects anomalies by utilizing the attribute correlations of nodes in each ego-network on the attribute network. In addition, residual analysis is another common method to measure the anomalies of nodes on an attribute network. In particular, Radar [10] characterizes the residuals of attribute information and their coherence with network information for anomaly detection. Further, ANOMALOUS [11] proposes joint learning of attribute selection and anomaly detection based on CUR decomposition and residual analysis. Despite the success of these methods on low-dimensional attribute graph data, they do not work well when graphs have complex structures and high-dimensional attributes [89, 90, 91, 92] due to the limitations of their shallow mechanisms.
@@
In contrast, unsupervised GAD approaches do not require any labels for model training. Similar to unsupervised AD for image data, unsupervised GAD approaches employ several unsupervised learning techniques to learn anomaly patterns on graph data, including data reconstruction [4, 34, 38], contrastive learning [15, 39, 40], and other auxiliary objectives [16, 17, 41, 42, 93]. For example, DOMINANT [4] is a reconstruction-based approach that employs a graph convolution autoencoder to reconstruct both the adjacency and attribute matrices simultaneously, assessing node abnormality through a weighted sum of the reconstruction error terms. Similarly, ComGA [34] is a community-aware attribute GAD framework based on tailored GCNs to capture local, global, and structural anomalies. CoLA [15], the first contrastive self-supervised learning framework for GAD, samples novel contrast instance pairs in an unsupervised manner, utilizing contrastive learning to capture local information. HCM-A [16] integrates both local and global contextual information, employs hop count prediction as a self-supervised task, and utilizes Bayesian learning to enhance anomaly identification. TAM [17] optimizes the proposed anomaly metric (affinity) on the truncated graph end-to-end, considering one-class homophily and local affinity. Nevertheless, all the above methods adhere to the conventional paradigm of “one model for one dataset”.
@@
Very recently, few studies apply ICL to graph learning and GNNs. As an initial attempt to use ICL for GNNs, PRODIGY [51] leverages a prompt graph-based framework to conduct few-shot ICL for node-level classification and edge-level prediction tasks. Further, UniLP [52] introduces ICL to resolve conflicting connectivity patterns caused by distributional differences between different graphs. Nevertheless, these methods require context samples in multiple classes for ICL during inference. Then, how to leverage one-class context samples with ICL in the scenario of generalist GAD remains an open question.
@@ Appendix D Discussion of Cross-Attentive In-Context Anomaly Scoring
D.1 Definitions of Single-Class and Multi-Class Normal
Dataset with single-class normal. In this type of dataset, the normal samples share the same pattern or characteristics. For example, in a network traffic monitoring system dataset, normal behavior might be defined by regular patterns of data packets exchanged between a specific set of IP addresses. Any deviation from this single, well-defined pattern, such as an unexpected spike in data volume or communication with unknown IP addresses, can be flagged as anomalous.
Dataset with multi-class normal. In this type of dataset, the normal samples are divided into multiple classes, each with distinct patterns or characteristics. For example, in a corporate email communication network dataset, normal data might be defined by regular patterns of email exchanges within specific departments, such as HR, IT, and Finance. Any deviation from these well-defined patterns, such as a sudden spike in emails between normally unconnected departments or an unusual volume of emails from an individual employee to external addresses, can be detected as anomalous.
D.2 ARC as One-Class Classification model
In this subsection, we first introduce the basic definition of one-class classification (OC) model, and then discuss the connection between one-class classification and cross-attentive in-context anomaly scoring module (C-AS for short).